# QueryIPI: Query-agnostic Indirect Prompt Injection on Coding Agents

### v1

@@ Abstract
Coding agents in modern IDEs orchestrate powerful tools and system-level actions, creating a high-stakes attack surface. Prior works on Indirect Prompt Injection (IPI) are mainly query-specific, rendering attacks unstable and with a relatively low success rate. We argue that a query-agnostic attack represents a far more severe threat to real-world LLM coding agents. However, it is fundamentally challenging, as the static query-agonistic payload must compete for the agent’s attention against diverse user query and other contextual prompts. Our key insight is that this challenge can be circumvented by exploiting a common vulnerability of LLM coding agents: the leakage of the agent’s internal prompt. Access to the prompt transforms the attack from an infeasible black-box search into a constrained, white-box optimization problem. Using this, we introduce QueryIPI, the first query-agnostic IPI method for realistic coding agents. QueryIPI uses the tool description as its malicious input channel and employs an iterative, prompt-based search to systematically refine this description. This method leverages the agent’s internal prompt in two key phases: initial seed generation, to create a description aligned with the agent’s role and security guardrails, and iterative reflection, to analyze failures and systematically rewrite the description to bypass those defenses. We evaluate QueryIPI on five simulated realistic coding agents, achieving an average success rate of 70%, 82%, and 87% with 2, 4, and 8 training samples, respectively, while the baseline achieves an average success rate of only 50%. Crucially, we demonstrate that the malicious tool descriptions generated in simulation successfully transfer to and compromise their real-world counterparts, proving that internal-prompt-aware, query-agnostic IPI is a practical and effective threat to contemporary coding agents.
@@
footnotetext: * Equal Contributionfootnotetext: Corresponding author (dongdong@cse.ust.hk)footnotetext: Source code: https://github.com/QueryIPI/QueryIPI.git
@@ 1 Introduction
Large language models (LLMs) powered coding agents have been widely deployed in major IDEs, including VS Code Copilot (cop, 2025), Cursor (Cursor, ), and Windsurf (win, 2025). Through protocols like the Model Context Protocol (MCP) (Anthropic, 2024), these agents evolved from simple code generators to powerful code development platforms that can interact with various external tools. This capability to interact with external tools brings coding agents a central and critical position in modern IDEs. With system-level operations and privileged context, coding agents can directly modify code and environments, creating a high-stake attack surface.
@@
Indirect prompt injection (IPI) has been shown to pose a significant security threat to LLM agents (Chen et al., 2024b; wunderwuzzi, 2025; Debenedetti et al., 2024; Greshake et al., 2023), especially the popular coding agents widely used in practice. Prior IPI research is mainly query-specific, as existing methods bind the attack to specific invoked tools in an ad-hoc fashion. One line of work embeds the attack payload in the invoked tool’s output (Debenedetti et al., 2024; Zhan et al., 2024; Wang et al., 2025b; Zhan et al., 2025). While another line of work embeds the attack payload within the tool’s description (Wang et al., 2025a; Labs, 2025a, b), where the existing attack paradigm is designed to trigger only when the user’s query directs the agent to use the specific invoked tool or a related category of functions. In both scenarios, the attack’s success hinges on the user’s query, rendering it unstable and with a relatively low success rate.
Figure 1: Comparison of Query-Specific and Query-Agnostic Indirect Prompt Injection. The left panel depicts a classic IPI attack, consistent with the threat models in recent benchmarks such as AgentDojo (Debenedetti et al., 2024) and Injecagent (Zhan et al., 2024). The attack is triggered only when a specific user query (①) causes the agent to invoke a compromised tool (②), which then injects malicious content (③). The probabilistic nature of this process is represented by dashed lines. In contrast, the right panel illustrates our query-agnostic attack, where the agent is compromised regardless of the user’s input (①), leading to a deterministic malicious outcome (②), as shown by the solid arrow. Malicious actions and content are highlighted in red.
We argue that a far more potent and realistic threat model is the query-agnostic attack. By definition, such an attack is designed to trigger reliably and execute its malicious payload regardless of the user’s specific input. As visualized in Figure 1, this transforms the threat from an unstable, low-probability event (left panel) into a stable and highly reproducible attack (right panel). But achieving this level of robustness is fundamentally challenging. The fundamental challenge in achieving robustness stems from the LLM’s attentional bias (Liu et al., 2023). For instance, attacks via tool returns are often effective because the model’s attention is naturally focused on processing the tool’s output. However, injecting through a tool description, for example, to establish a connection with a malicious MCP server, is substantially more difficult. Experimental evidence confirms this, showing that prompts designed for tool-return IPI attacks have a near-zero success rate in a tool-description setting (Wang et al., 2025a). When injecting via a tool description, the injected content faces significant interference from the user query, the system prompt (including embedded defenses (Liu et al., 2024; Chen et al., 2024a; Hines et al., 2024; san, 2023; ins, 2023)), and other competing tools. As a query-agnostic attack, the injected content cannot leverage the user’s query to direct the model’s focus, making it exceptionally challenging to overcome this divided attention.
We observe that this challenge can be circumvented by a common vulnerability: the leakage of the agent’s internal prompt. These prompts are often exposed through various channels, including direct prompt injection surfaces (Xie et al., 2025; x1xhlol, 2025) or via white-box access to open-source agents (Team, 2025; Chen et al., 2021). Once the internal prompt is known, an attacker can analyze the agent’s internal instructions, including its operational logic and prescribed safety guardrails. This knowledge turns the black-box search for an effective attack into a more constrained, white-box optimization problem. Instead of blindly guessing, the attacker can craft a payload that targets explicit weaknesses in the agent’s logic, thereby reducing the search space and increasing the stability and transferability of the attack.
@@
We present QueryIPI, the first query-agnostic IPI method designed for realistic coding agents. To achieve this query-agnostic behavior, QueryIPI leverages the tool description as its malicious input channel. Our approach employs an iterative, prompt-based search method to systematically refine this malicious description until it reliably triggers the desired payload. This method critically leverages the agent’s internal prompt in two key phases: initial seed generation and iterative reflection. Specifically, during initial seed generation, QueryIPI conditions the generation process on the internal prompt to create a seed description that aligns with the agent’s prescribed role, tool-use patterns, and embedded security guardrails. Then, during the reflection phase of each iteration, it explicitly uses the internal prompt to analyze failures and rewrite the description, enabling it to craft reformulations that systematically target and bypass the agent’s internal defenses.
@@
To automate our experiments, we simulated five realistic coding agents: Cursor (Cursor, ), Windsurf (win, 2025), Cline (Team, 2025), Copilot (cop, 2025), and Trae (tra, 2025). Because our attack does not depend on the output of a tool’s invocation, our evaluation can focus solely on the agent’s initial response to a user query, without requiring actual tool interaction. Our method proves highly effective. On five simulated realistic coding agents, QueryIPI achieves an average success rate of 70%, 82%, and 87% with 2, 4, and 8 training samples, respectively, while the baseline achieves an average success rate of only 50%. Crucially, we also demonstrate that the malicious tool descriptions generated through our simulated environment successfully transfer to and compromise the actual, real-world versions of these coding agents.
@@
  • We are the first to define and formalize the query-agnostic IPI attack, a novel threat model for LLM agents.
  • We propose QueryIPI, an automated method that leverages the agent’s internal prompt to systematically generate these attacks.
@@
2 Related Work – IPI Attack
Indirect Prompt Injection is a significant security threat where an LLM is compromised by malicious instructions embedded in external data sources it consumes (Greshake et al., 2023). While this threat applies broadly to various LLM-integrated applications, our work concentrates on the more complex and autonomous paradigm of LLM agents.
@@ 2 Related Work – IPI Attack
The most common line of work explores IPI against agents through malicious content embedded in tool outputs or retrieved documents. Several benchmarks have been developed specifically to evaluate LLM agents in this context. AgentDojo (Debenedetti et al., 2024) and Injecagent (Zhan et al., 2024) provide environments where an agent’s tools can return poisoned results. The vwa-adv benchmark (Wu et al., 2024) extends this agent-centric paradigm to multimodal scenarios. Building on these environments, various attack strategies against agents have been proposed. AGENTVIGIL (Wang et al., 2025b) introduced a generic black-box red-teaming method, but its success still relies on the agent selecting the compromised tool. Other works have specialized this attack vector for specific agent domains, such as web agents that process poisoned content (AdvAgent (Xu et al., 2024)) or agents that handle structured data (StruPhantom (Feng and Pan, 2025)). Adaptive attacks against agents (Zhan et al., 2025) also fits this model. A common thread unites all these approaches: the user must issue a query that compels the agent to invoke the specific tool or retrieve the document containing the payload, thus rendering the attack opportunistic.
A second line of work injects malicious content into an agent’s tool descriptions, for example, to establish a connection with a malicious MCP server, or other metadata (Wang et al., 2025a; Labs, 2025a, b). This vector presents a distinct challenge, as research shows that naively porting attack techniques from the tool-output setting yields very low success rates(Wang et al., 2025a). The threat model in these works is query-specific: the attack paradigms are designed to trigger only when the user’s query directs the agent to use the specific compromised tool or a related category of functions. A recent work in this area is ToolHijacker (Shi et al., 2025), which also crafts malicious tool descriptions. However, this retrieval-based tool selection mechanism is not applicable to realistic coding agent settings.
3 Preliminary
Threat Model of Query-Agnostic IPI Attack. We model an agent, $\mathcal{A}$, as a system defined by its system prompt, $\mathcal{P}_{\text{agent}}$, and a toolset, $\mathcal{T}$, with a specific instance denoted as $\mathcal{A}_{\mathcal{P}_{\text{agent}},\mathcal{T}}$. In a benign scenario, this agent takes a user query, $\mathcal{Q}$, as input to produce an intended behavior, $\mathcal{B}^{B}$, and a corresponding response, $\mathcal{R}^{B}$, which we express as $\mathcal{A}_{\mathcal{P}_{\text{agent}},\mathcal{T}}(\mathcal{Q})=(\mathcal{B}^{B},\mathcal{R}^{B})$. Our threat model considers a query-agnostic attack where an attacker, who has full knowledge of internal prompt, including $\mathcal{P}_{\text{agent}}$ and $\mathcal{T}$, and can observe the agent’s outputs $(\mathcal{B},\mathcal{R})$, crafts and injects a malicious tool, $t_{mal}$, into the agent’s environment. This action creates a compromised toolset, $\mathcal{T}^{\prime}=\mathcal{T}\cup{t_{mal}}$, effectively altering the agent’s instance. The attack is successful if, for any arbitrary user query $\mathcal{Q}$, this compromised agent, $\mathcal{A}_{\mathcal{P}_{\text{agent}},\mathcal{T}^{\prime}}$, is manipulated into performing a predetermined malicious action, $a_{\text{mal}}$, as part of its behavior $\mathcal{B}^{T}$. This malicious outcome is represented by the equation $\mathcal{A}_{\mathcal{P}_{\text{agent}},\mathcal{T}^{\prime}}(\mathcal{Q})=(\mathcal{B}^{T},\mathcal{R}^{T})$, where the execution of the payload is triggered regardless of the user’s query. We will discuss the scenario where the attacker does not have complete knowledge of the internal prompt in the Section˜5.3.
Formulation. Our objective is to automatically generate a malicious tool description, $d$, that reliably triggers the predetermined malicious action $a_{\text{mal}}$. To make the optimization practical, we formulate it over a fixed training set of diverse user queries, denoted as ${\mathcal{Q}_{k}}_{k=1}^{K}$. We aim to find the optimal description $d_{a_{\text{mal}}}^{*}$ that maximizes the cumulative success score across all queries in this set:
$d_{a_{\text{mal}}}^{*}=\arg\max_{d\in\mathcal{D}}\sum_{k=1}^{K}s(\mathcal{Q}_{k},\mathcal{A}_{\mathcal{P}{\text{agent}},\mathcal{T}^{\prime}(d)}(\mathcal{Q}_{k}),a_{\text{mal}})$ (1)
Here, $K$ is the total number of queries in the training set, and the summation $\sum_{k=1}^{K}$ represents the aggregation of success scores over this entire set. $\mathcal{D}$ is the search space for the tool description $d$. We expressly define $\mathcal{D}$ in Appendix˜A. $\mathcal{T}^{\prime}(d)=\mathcal{T}\cup{t_{mal}(d)}$ is the agent’s toolset augmented with the malicious tool. The scoring function $s(\cdot,a_{\text{mal}})$ evaluates the agent’s performance for each individual query $\mathcal{Q}_{k}$ by checking for the successful execution of the specific target action $a_{\text{mal}}$. We will describe this scoring function in detail in Section˜4.2.
@@ 4 QueryIPI
To solve the optimization problem defined in the previous section and find the optimal malicious description $d_{a_{\text{mal}}}^{*}$, we employ an iterative optimization procedure, which is formally outlined in Algorithm 1. As shown in the algorithm, the process is initialized with a single random string as the initial seed. It then enters the main iterative loop, which runs for a predetermined number of generations, $G$. In each generation, a Mutation function generates a set of new variants from the current best-performing description (Section˜4.1). These variants are then evaluated and ranked by our Scoring function (Section˜4.2). The top-scoring description is selected to proceed to the next generation. After completing all generations, the algorithm returns the final candidate description as the optimal result.
Algorithm 1 Iterative Optimization Procedure
2:Output: Optimal malicious description.
3:// Use a random string as the inital seed
[…]
5:for $g=1$ to $G$ do
6:  $D_{\text{var}}\leftarrow\text{Mutation}(d_{\text{cand}},N)$ // Section˜4.1
7:  for each description $d$ in $D_{\text{var}}$ do
8:   $s_{d}\leftarrow Scoring(d)$ // Section˜4.2
9:   // Early stop mechanism
10:   if $s_{d}==FullScore$ then return $d$
11:   end if
12:   $Scores[d]\leftarrow s_{d}$
13:  end for
14:  // Choose the top score
15:  $(d_{\text{cand}},s_{\text{cand}})\leftarrow\text{SelectTop}(Scores)$
16:  $Scores\leftarrow\{\}$
17:  $Scores[d_{\text{cand}}]\leftarrow s_{\text{cand}}$
18:end for
19:
20:return $d_{\text{cand}}$
4.1 Mutation Function

### v3

@@ Abstract
Coding agents in modern IDEs orchestrate powerful tool and high-privilege system access, creating a high-stakes attack surface. Prior work on Indirect Prompt Injection (IPI) is mainly query-specific, i.e., it requires a specific user query as trigger to detonate the attack, leading to poor generalizability across diverse attack scenarios and tasks. We propose a new attack paradigm: query-agnostic IPI that reliably triggers and executes malicious payload under an arbitrary user query, posing a severe threat to real-world coding agents.
The query-agnostic IPI requires reliable instruction-following of the malicious payload regardless of the user query content. Our key insight is that the malicious payload should leverage the invariant prompt context of coding agent rather than the variant user query. We identify the internal prompt (i.e., system prompt and internal tool description) as the system invariant of the coding agent. We then use the system invariant to guide the generation of query-agnostic IPI payload via an effective blackbox optimization.
We present QueryIPI, an automated framework to launch a query-agnostic IPI for coding agent. QueryIPI uses the tool description as optimizable payload and employs an iterative, prompt-based search to refine it. QueryIPI leverages system invariant of coding agent in two key phases: 1) initial seed generation, to create a description aligned with agent’s native style and tool conventions; 2) iterative reflection, to analyze failures and systematically rewrite the description, aiming to resolve instruction following failure and safety refusal.
We evaluate QueryIPI on five simulated real-world coding agents, achieving an average success rate of 70%, 82%, and 87% with 2, 4, and 8 training samples, respectively, while the best-performing baseline achieves an average success rate of only 50%. Crucially, we demonstrate that the malicious tool descriptions generated in simulation successfully transfer to and compromise the real-world coding agent, highlighting that query-agnostic IPI is highly effective in practice.
@@
footnotetext: * Equal Contributionfootnotetext: Corresponding author (pchen@fudan.edu.cn)footnotetext: Source code: https://github.com/QueryIPI/QueryIPI.git
@@ 1 Introduction
Large language models (LLMs) powered coding agents have been widely deployed in mainstream coding Integrated Development Environment (IDEs), including VS Code Copilot (cop, 2025), Cursor (Cursor, ), and Windsurf (win, 2025). Through protocols such as the Model Context Protocol (MCP) (Anthropic, 2024), these agents evolved from simple code completion/generation tools to powerful platforms capable of interacting with various external tools. This capability elevates coding agents to a central and critical position in modern IDEs. With high-level system access (e.g., file system access and shell command privilege), coding agents can influence the entire system, creating a high-stakes attack surface.
Figure 1: Comparison of Query-Specific and Query-Agnostic Indirect Prompt Injection. The left panel depicts a query-specific IPI where attack success is conditioned on the user query: benign queries remain secure (green shields), and the attack is only triggered by a specific query (red warning). Consequently, the compromise path is conditional, represented by the dashed red arrow. In contrast, the right panel illustrates our query-agnostic IPI. The agent is consistently compromised under an arbitrary user query. Arbitrary user query can trigger the malicious payload, resulting in a reliable attack, as shown by the solid red arrow.
@@
Indirect prompt injection (IPI) poses a significant security threat to LLM agents (Chen et al., 2024; wunderwuzzi, 2025; Debenedetti et al., 2024; Greshake et al., 2023), especially widely-used coding agents. Prior IPI work has mainly been in a query-specific setting, in which attacks require a specific user query as trigger to detonate. One line of work embeds the malicious payload in the output of the invoked tool (Debenedetti et al., 2024; Zhan et al., 2024; Wang et al., 2025b; Zhan et al., 2025). While another line of work embeds the malicious payload within the tool’s description (Wang et al., 2025a; Labs, 2025a, b), where the existing attacks can be triggered only when a specific tool is invoked by user query. In both cases, the attack’s success hinges on specific user queries as attack triggers, leading to poor generalizability to diverse attack scenarios and tasks. In practice, it is unreliable and has a relatively low success rate.
We introduce a far more potent and realistic attack paradigm: query-agnostic IPI. It reliably triggers and executes the malicious payload under arbitrary user queries. As visualized in Figure 1, this transforms the paradigm from a conditional query-specific attack (left) to an unconditional query-agnostic attack (right).
However, achieving query-agnostic IPI in the real-world coding agent is challenging, primarily due to the huge and diverse prompt context. The powerful and complex capabilities of coding agents, such as complex tool invocation, lead to a huge prompt context where the malicious payload injected by the attacker is easily ignored by the LLM. Meanwhile, the semantic misalignment between the arbitrary user query and the malicious payload further contributes to the poor instruction-following of the malicious payload.
To tackle this challenge, our key insight is that the malicious payload should leverage the invariant prompt context of the coding agent rather than the variant user query. After careful investigation of the coding agent’s prompt structure, we identify the internal prompt (i.e., system instructions and internal tool descriptions) as a system invariant that is always digested by LLM and uniformly affects the coding agent’s behavior across user query content. We then leverage the system invariant of the coding agent to guide the generation of query-agnostic IPI via an effective black-box optimization. Additionally, we show that obtaining these system invariants of real-world coding agents is highly viable and practical. These internal prompts are frequently exposed via open-source frameworks Team (2025); Chen et al. (2021) or highly effective extraction attacks (Xie et al., 2025; Peng et al., 2025; Agarwal et al., 2024). Driven by commercial incentives, these prompts are also actively targeted and leaked in public repositories (x1xhlol, 2025). Although several defense prototypes against prompt leakage have been proposed (Cao et al., 2025; Pape et al., 2025), they are far from real-world deployment due to the potential performance degradation.
@@
We present QueryIPI, the first query-agnostic IPI attack for real-world coding agents. QueryIPI uses the external tool description as its malicious payload. Our approach employs an iterative, prompt-based search method to systematically refine this malicious tool description until it reliably triggers the desired adversarial behavior of the coding agent. This method leverages coding agent’s system invariant (i.e., internal prompt) in two key phases: initial seed generation and iterative reflection. During initial seed generation, QueryIPI generates a seed description that aligns with the agent’s prescribed role, tool-use patterns, and embedded security guardrails, using internal prompt information. Then, in the iterative reflection phase, it explicitly uses the internal prompt to analyze failures and rewrite the description, enabling it to craft reformulations that systematically target and bypass the agent’s internal defenses.
@@
To automate our experiments, we simulate five realistic coding agents: Cursor (Cursor, ), Windsurf (win, 2025), Cline (Team, 2025), Copilot (cop, 2025) and Trae (tra, 2025). Because our attack does not depend on the output of a tool’s invocation, our evaluation can focus solely on the agent’s initial response to a user query, without any interaction with the tool. Our evaluation shows that QueryIPI is highly effective. Across five simulated real-world coding agents, QueryIPI achieves average success rates of 70%, 82%, and 87% with 2, 4, and 8 training samples, respectively, whereas the best-performing baseline achieves only 50%. Crucially, we show that malicious tool descriptions crafted in our simulated environment retain their efficacy when deployed against real-world coding agents, successfully compromising them.
@@
  • We pinpoint the limitation of query-specific IPI attack: reliance on a specific user query to detonate.
  • We give the first formalization of the query-agnostic IPI attack for the coding agent.
  • We identify the internal prompt as system invariant of coding agents. We then propose an automated framework QueryIPI to generate query-agnostic IPI for coding agent with system invariant.
@@
2 Query-Specific IPI
To the best of our knowledge, prior works of Indirect Prompt Injection (IPI) (Greshake et al., 2023) on coding agent are all query-specific as they require a particular trigger (a specific user query like "write a maze game" as shown in Fig 1) to detonate the attack.
@@ 2 Query-Specific IPI
One line of work explores IPI by embedding malicious content in tool outputs or in retrieved documents. AgentDojo (Debenedetti et al., 2024) and Injecagent (Zhan et al., 2024) show that tool invocation of an LLM agent can return poisoned and dangerous content on a particular user query. The vwa-adv benchmark (Wu et al., 2024) extends this query-specific IPI paradigm to multimodal domain. AGENTVIGIL (Wang et al., 2025b) introduced a generic black-box red-teaming method, but its success still relies on a specific compromised tool invocation. Other works (Xu et al., 2024; Feng and Pan, 2025; Zhan et al., 2025) have specialized this attack to various agent domains.
Another line of work focuses on embedding malicious content in attacker-controlled tools (tool description or metadata) by constructing benign-looking or spoofed packages (Hou et al., 2025; Li and Gao, 2025; securelist, 2025; Christian Posta, 2025). For instance, attackers can manipulate tool definitions to covertly establish connections with malicious MCP servers (Wang et al., 2025a; Labs, 2025a, b). They require a specified user query to either invoke an attacker-controlled tool (Labs, 2025a) or a benign tool abused by attackers (Labs, 2025b). Similarly, recent works such as MCPTox (Wang et al., 2025a) and ToolHijacker (Shi et al., 2025) are limited to scenarios where the user performs a specific task category.
Limitations. These approaches share a fundamental limitation: reliance on a specific user query. The carefully crafted malicious payloads that influence LLM agent behavior only when the user sends a particular user query, and have no impact on the LLM agent under any other non-trigger user query. As a result, query-specific IPIs are often restricted to a few specified attack scenarios and tasks and fail to generalize to diverse application scenarios.
3 New Paradigm: Query-Agnostic IPI
To overcome above limitations of query-specific IPI, we introduce a new attack paradigm for coding agents: Query-Agnostic IPI. Unlike prior methods that depend on a specific user query as an attack trigger, a query-agnostic IPI is designed to hijack the coding agent’s behavior under arbitrary user query. This paradigm shift transforms the IPI attack from a conditional query-specific setting to an unconditional query-agnostic one. A query-agnostic IPI ensures that the malicious payload is always executed by the coding agent regardless of the user query content. Therefore, our approach poses a significantly severe and practical threat to real-world coding agents by generalizing to arbitrary user queries and task scenarios.
Threat Model. We define a coding agent $\mathcal{A}$ includes system prompt ${p}_{{sys}}$, user query q{q} and a toolset $\mathcal{T}$(a set of internal tools and external tools). Its intended behavior is denoted as b{b}. We then have the model of coding agent as: $\mathcal{A}_{{p}_{{sys}}}^{\mathcal{T}}({q})=b$. We assume attacker has no access to any runtime state of victim coding agent, but he or she knows the specific version number of victim coding agent. Hence, an attacker can set up a local copy of the same version of the victim coding agent to test the exploit, since a coding agent can be easily obtained from the internet. The attacker crafts a malicious tool $t_{mal}$ and tricks the victim coding agent $\mathcal{A}$ into integrating $t_{mal}$ into its toolset $\mathcal{T}$ via package, name spoofing or any other supply chain attacks (Hou et al., 2025; Li and Gao, 2025; securelist, 2025; Christian Posta, 2025). The compromised toolset is then denoted as: $\mathcal{T}^{\prime}=\mathcal{T}\cup\{t_{mal}\}$. The attack goal is to hijack the behavior of coding agent $\mathcal{A}$ under arbitrary user query $q$ such that it always output a targeted malicious behavior $b_{{mal}}$, represented as $\mathcal{A}_{{p}_{{sys}}}^{\mathcal{T^{\prime}}}({q})\equiv b_{mal},\forall q\in Q$, where $Q$ denotes the set of all possible user queries.
@@ 4 QueryIPI
We introduce QueryIPI, an automated framework that can generate malicious tool description for launching query-agnostic IPI attack on the coding agent. We consider the internal prompt of the coding agent (e.g., system prompt and tool descriptions of internal tools) as system invariants that uniformly impact the LLM agent’s behavior regardless of the user query content. We then use system invariants to guide the generation of query-agnostic IPI payload through an effective blackbox optimization.
4.1 Internal Prompt as System Invariant
Motivation. In query-agnostic IPI, the malicious payload should elicit reliable adversarial instruction-following in the coding agent under an arbitrary user query. However, it is quite challenging due to the huge and diverse prompt context of a real-world coding agent.
The prompt context primarily consists of system prompt, user query and tool-invocation content (i.e., tool descriptions). Specifically, to support the complex and powerful capabilities of coding agents, this context is heavily burdened by detailed descriptions of tool usage, intricate planning logic embedded in the system prompt, and various prompt-level security constraints, all of which collectively result in an exceptionally large. As a result, the malicious payload is easily lost in the huge prompt context, where the “lost in the middle” phenomenon (Liu et al., 2023; Li et al., 2024) often causes LLM to ignore any instruction in the malicious payload. Moreover, the adversarial instruction-following heavily relies on the semantic alignment of user query and malicious payload. But it’s impossible to align the semantics of malicious payload with an arbitrary user query.
System Invariant. To achieve reliable adversarial instruction-following, we carefully investigate the prompt context of coding agent and discover a crucial “system invariant” amidst the dynamic interactions: the internal prompt, which comprises the system instructions and descriptions of the internal toolset. Unlike the variable user query, this system invariant is consistently digested by the LLM during every inference to guide role-playing, safety enforcement, and next-step planning, regardless of user query and historical context. Hence, it naturally fits our query-agnostic setting. By strategically aligning our malicious instructions with this system invariant, we can enable a stable adversarial instruction-following and further achieve a query-agnostic attack.
Feasibility. We argue that obtaining internal prompts of coding agents is highly practical and viable. First, many coding agents, such as Cline (Team, 2025), are open-source, providing direct access to their system instructions. Second, regarding internal prompts of proprietary LLM models, recent research demonstrates that prompt extraction attacks are highly effective in practice. RepeatLeakage (Peng et al., 2025) reports failure rates below 10%, while other recent studies achieve success rates reaching 99.9% on GPT-4 via multi-turn strategies (Agarwal et al., 2024). Third, the high commercial value of these coding agents incentivizes adversaries to actively steal their internal prompts, leading to the creation of open-source repositories (x1xhlol, 2025) that collect and expose leaked system prompts. Lastly, although several defense prototypes have been introduced in academia (Cao et al., 2025; Pape et al., 2025), they are barely deployed in real-world production environments due to potential performance degradation. We will discuss the implications of using approximated prompts versus exact internal prompts in Section˜5.3.
4.2 System Invariant-Guided Optimization
Given the system invariant of coding agent as feedback, we formulate the generation of malicious payload as a blackbox optimization. We define the tool description of a malicious external tool as the optimizable payload. The optimization goal is to generate a malicious tool description $d_{{mal}}$ such that it can reliably triggers targeted malicious behavior $b_{{mal}}$ over a dataset of $k$ diverse and representative user queries, denoted as ${Q}=\{q_{i}|i=1,2,...,k\}$.To construct Q{Q}, we randomly sample queries from a large-scale real-world conversation dataset. Detailed settings are provided in Section˜5. Concretely, we aim to find the optimal description $d_{{mal}}^{*}$ that maximizes cumulative success score across all queries in $Q$:
$d_{mal}^{*}=\operatorname*{arg\,max}_{d_{mal}\in{D}}\sum_{i=1}^{k}\mathds{1}\left(\mathcal{A}_{p_{sys}}^{\mathcal{T}^{\prime}}(q_{i})=b_{mal}\right)$ (1)
The summation represents the aggregated success scores over the dataset $Q$. D{D} denotes all possible malicious tool descriptions. We expressly define $\mathcal{D}$ in Appendix˜C. Note that $d_{mal}$ is included in the compromised toolset $\mathcal{T}^{\prime}$. The function $\mathds{1}$ is a binary indicator that evaluates the success of the attack. It assigns 1 if the specific target action $a_{mal}$ is successfully executed, and 0 otherwise.
To solve this optimization problem, we employ an iterative evolutionary procedure (formally outlined in Algorithm 1 in Appendix˜A). The process initializes by utilizing the Mutation function to generate an initial set of seed descriptions. It then iterates through $G$ generations. In each generation, the current set of candidates is first evaluated by the Scoring function. Based on these scores, the best-performing description is selected. Subsequently, the Mutation function generates a new set of candidates for the next iteration by leveraging this top-performing seed. This cycle repeats until the maximum number of generations is reached or a full success score is achieved.